# Different physical and numerical sources of scatter in the $M_{\star}$-$M_{\mathrm{BH}}$ relation and their connection to galaxy evolution

### v1

@@ 3.3 The scatter in TNG from different origins
Based on the description in Section 3.1, the sources of the scatter in the $M_{\rm BH}-M_{\star}$ relation can be broadly divided into numerical and physical ones. As described in the previous section, with the help of two sets of toy simulations we can obtain estimates for the numerical contribution to the scatter, and separate the intrinsic scatter into contributions from hierarchical merging and BH growth by gas accretion. These results are summarized in the upper panel of Figure 7. Note that all of the results here are obtained from galaxies fulfilling the selection criteria discussed in Section 2.2.1. The dash-dotted and solid red lines show the scatter and intrinsic scatter in TNG100 cosmological simulations after subtracting the scatter from numerical origins. The difference between scatter and intrinsic scatter is $\sim$ 0.02, and our estimates for $\beta$, $\sigma_{\rm BH}$, and $\sigma_{\star}$ are $\sim$ 1, 0.01 and 0.01, respectively. The physical sources of scatter from hierarchical merging and BH accretion are also shown in the upper panel of Figure 7.
Figure 7: Upper panel: The time evolution of the scatter of the $M_{\rm BH}-M_{\star}$ relation, the intrinsic scatter after correcting for the scatter from numerical origins, the scatter due to hierarchical merging, and the the scatter due to BH accretion in the TNG100 simulation. Bottom panel: The time evolution of fractional contributions to the scatter due to numerical origins, hierarchical merging, and BH accretion in the TNG100 simulation.
Note that the scatter from hierarchical merging is relatively low and increases with decreasing redshift, which arises because there massive galaxies increasingly merge with small galaxies that do not strictly lie on the $M_{\rm BH}-M_{\star}$ relation. In contrast, the inferred scatter from BH accretion decreases with decreasing redshift, and becomes consistent with zero at $z\lesssim 0.5$. This can be easily understood since the massive galaxies will tend to quench with decreasing redshift due to the AGN feedback, so that the contribution of BH accretion to the scatter scatter becomes lower and lower with time for massive galaxies and disappears at low redshift.
@@
From the figure, it can also be seen that the scatter from the BH accretion decreases from a dominating $80\%$ at $z=3$ to something consistent with $\sim 0\%$ at $z\sim 1$. In contrast, the relative contribution of hierarchical merging increases from being negligible at $z=3$ to $\sim 40\%$ at $z\sim 0$, indicating that hierarchical merging contributes significantly to the scatter of the $M_{\rm BH}-M_{\star}$ relation at $z=0$ in the TNG simulations. It is however interesting to note that numerical sources are dominating the scatter at $z<2$ in the TNG model, therefore it would be misleading to directly compare the scatter obtained from the simulations with observations without a corresponding correction. In fact, the scatter seen in TNG at low redshift would be noticeable lower in the absence of numerical noise.
Figure 8: Upper panel: The time evolution of the intrinsic scatter of the $M_{\rm BH}-M_{\star}$ relation in TNG100, Illustris, EAGLE, and MTNG. Bottom panel: The time evolution of the fraction of the scatter due to BH accretion with respect to the total scatter in the TNG100, Illustris, EAGLE, and MTNG simulations, based on the assumption that the scatter from hierarchical merging is similar in all simulations.
Figure 9: Upper panel: The time evolution of the total and intrinsic scatter in the $M_{\rm BH}-M_{\star}$ relation in the TNG100 and SeedVar simulations. Middle panel: The time evolution of the scatter of the BH masses in different BH mass bins, and for different redshifts, as labelled. Lower panel: The time evolution of fractional contributions of different sources of scatter in the BH masses at different redshifts.
@@ 3.4 Other cosmological simulations and observational comparison
We now try to apply the lessons learned from TNG about the different sources of scatter in the $M_{\rm BH}-M_{\star}$ relation to the Illustris, EAGLE, and MTNG runs. Instead of carrying out similar numerical experiments as described in Section 3.2, we simply assume that the scatter due to hierarchical merging is similar in these simulations compared to TNG. This should certainly be a good assumption for the MTNG simulation, which shares the simulation code and an almost identical galaxy formation model with TNG100. But it is somewhat less obvious for the Illustris and EAGLE simulations. Note, however, that their BH seeding scheme is conceptionally very similar to TNG, suggesting that the time evolution of the scatter due to hierarchical merging should be broadly similar.
@@
However, because the BH feedback and galaxy formation models in Illustris and EAGLE are quite different from TNG, the scatter due to numerical sources may be different. Thus we cannot directly apply the numerical scatter we obtained from TNG to the EAGLE and Illustris simulation data. Fortunately, Borrow et al. (2023b) measured the scatter due to the numerical variations in the EAGLE simulation. Based on their work, the variance of the BH mass $\sigma_{\rm BH}^{2}$ due to numerical origins in massive galaxies is around $0.01-0.03$, and for the stellar mass the scatter $\sigma_{\star}^{2}$ in massive galaxies is around 0.01 from numerical origins, which is comparable to what we inferred for TNG. Unfortunately, for Illustris there is no work in the literature that has reported the scatter from numerical origins. As Illustris shares many aspects of the galaxy formation model with TNG, and also uses the same code, we however expect these two simulations to be similar in practice with respect to the numerical scatter, so we shall assume the TNG numerical scatter also for Illustris.
@@
Based on above assumptions, we can obtain the intrinsic scatter and the fractional contributions of different sources of scatter in these four cosmological simulations. We show the results in Figure 8, where the upper panel displays the intrinsic scatter in the four simulations, while the black dash line at $z\sim 0$ is the intrinsic scatter obtained by Kormendy & Ho (2013), who report a value of $\sim 0.28$ dex. It can be seen that TNG100 and MTNG have very similar intrinsic scatter, as expected, which amounts to $\sim 0.1$ dex at $z\sim 0$. In contrast, Illustris and EAGLE show similar scatter, with values comparable to the observations. Since hierarchical merging reduces the scatter (Peng, 2007; Hirschmann et al., 2010; Jahnke & Macciò, 2011), and a tighter co-evolution of BHs and their host galaxies also regulates systems to be close to the $M_{\rm BH}-M_{\star}$ relation (Kormendy & Ho, 2013; Heckman & Best, 2014; Sun et al., 2015; Zhuang & Ho, 2023), we conclude that the low level of scatter in TNG100 and MTNG reflects a more efficient quenching of massive galaxies and a particularly tight self-regulation of the $M_{\rm BH}-M_{\star}$ relation in this model.
@@
The bottom panel of Figure 8 shows the fractional contribution of the variance from BH accretion to the total variance. This contribution by BH accretion is $>50\%$ for TNG100 and MTNG at $z=3$ and then becomes negligible at low redshift. In contrast, in Illustris and EAGLE, the fraction of the variance from BH accretion is $\sim 80\%$ across the redshift range $z=3$ to $z=0$. This result indicates that the $M_{\rm BH}-M_{\star}$ relation is mainly regulated by hierarchical merging in TNG100 and MTNG at low redshift, with BH accretion becoming subdominant there, whereas Illustris and EAGLE appear to be less strongly quenched, so that variations in the BH mass by gas accretion are still more important. We will further discuss this issue in Section 4.2.
@@ 3.5 Scatter due to BH seed mass variations
As mentioned in Section 3.1, the scatter in the $M_{\rm BH}-M_{\star}$ relation predicted by the simulations may also be affected by variations of the BH seed mass. The seed mechanism itself is presently unknown, and many different possibilities are actively discussed, including the run-away collapse of dense star clusters, the collapse of massive gas clouds, or the remnants of unusually massive pop-III stars. While the latter would still be comparatively light, the former may perhaps produce seeds with masses in excess of $10^{5}\,{\rm M}_{\odot}$. The seed masses could exhibit substantial variations (e.g. Bhowmick et al., 2024b), and especially if they are relatively massive the initial spread may still affect the scatter of the BH masses later on.
@@
Providing a more elaborate physical model for BH seeding is beyond the scope of the current work. Instead we carry out a numerical experiment where we investigate the impact of varying the BH seed mass by running an extra cosmological simulation where we impose a random scatter of $0.5$ dex on the BH seed mass introduced in a small TNG-like simulation (hereafter SeedVar) with box size $25\,h^{-1}{\rm cMpc}$ and $2\times 256^{3}$ resolution elements.
@@
Figure 10: The scatter in the BH mass of galaxies of different stellar mass at different redshifts in the TNG100, Illustris, EAGLE, and MTNG simulations. The black line represents the scatters of BH mass for all galaxies. The blue line gives the scatter only for blue galaxies, while the red line correspondingly is for red galaxies. For the TNG100, Illustris, and MTNG simulations, the blue galaxies are defines as galaxies with color index $(g-r)<0.6$ while the red galaxies are the galaxies with color $(g-r)>0.7$. For EAGLE simulation, the blue galaxies are defined as galaxies with color $(g-r)<0.75$ while the red galaxies are those with color $(g-r)>0.75$. Upper-left panel: TNG100 simulation. Upper-right panel: Illustris simulation. Lower-left panel: EAGLE simulation. Lower-right panel: MTNG simulation.
@@
Figure 11: Left panels: The median value of the specific SFR as a function of stellar mass at different redshifts in the TNG100, Illustris, EAGLE, and MTNG simulations. Right panels: The median value of the BHAR as a function of stellar mass at different redshifts in TNG100, Illustris, EAGLE, and MTNG.
@@
The result is presented in Figure 9. The upper panel displays the evolution of variance and intrinsic scatter of the BH mass for the massive galaxies in the TNG100 and SeedVar simulations. The intrinsic variance is calculated by assuming that the numerical scatter in SeedVar is similar to that in TNG100. Comparing the resulting intrinsic variance in the SeedVar and TNG100 simulations, the scatter is higher in SeedVar than in TNG100 except for $z\sim 0$, where the intrinsic variance in SeedVar is comparable to that in TNG100. The variance in the SeedVar simulation at $z\lesssim 0.5$ experiences a steep decrease, from $\sim 0.1$ to $\sim 0.01$, i.e. the intrinsic scatter drops from $\sim 0.3$ dex to $\sim 0.1$ dex. This result is likely due to the effective galaxy quenching in TNG simulations.
@@
We further investigate the scatter of BH masses in different BH mass bins and at different redshifts in the middle panel of Figure 9. The BHs in SeedVar show high scatter in low-mass BH bins due to the large variation of BH seeds. However, this increase in the scatter quickly decays away at $M_{\rm BH}\sim 10^{8}\,{\rm M_{\odot}}$, which can be interpreted as a loss of memory of the BH seed variations, and becomes $\lesssim 0.1\,{\rm dex}$ for massive BHs. This critical BH mass is the mass at which the impact of AGN feedback becomes significant, and the galaxies begin to be quenched (Zinger et al., 2020). Since the quenched galaxies will have relatively low accretion rate due to a lack of cold gas, and the growth of BHs in quenched galaxies is mainly occurring due to mergers, we can infer that the quick drop of the scatter of the BH mass is due to galaxy quenching at $M_{\rm BH}\sim 10^{8}\,{\rm M_{\odot}}$.
@@
Figure 12: Upper panel: The difference of the sSFR for the stellar mass bins in the ranges $10^{10-11}\,{\rm M_{\odot}}$ and $10^{10.5-11.5}\,{\rm M_{\odot}}$ in TNG100, Illustris, EAGLE, and MTNG. Lower panel: The difference of the BHAR in units if the Eddington rate for the stellar mass bins in the ranges $10^{10-11}\,{\rm M_{\odot}}$ and $10^{10.5-11.5}\,{\rm M_{\odot}}$ in TNG100, Illustris, EAGLE, and MTNG.

### v2

@@ 3.3 The scatter in TNG from different origins
Based on the description in Section 3.1, the sources of the scatter in the $M_{\rm BH}-M_{\star}$ relation can be broadly divided into numerical and physical ones. As described in the previous section, with the help of two sets of toy simulations we can obtain estimates for the numerical contribution to the scatter, and separate the intrinsic scatter into contributions from hierarchical merging and BH growth by gas accretion. These results are summarized in the upper panel of Figure 7. Note that all of the results here are obtained from galaxies fulfilling the selection criteria discussed in Section 2.2.1. The dash-dotted and solid red lines show the scatter and intrinsic scatter in TNG100 cosmological simulations after subtracting the scatter from numerical origins. The difference between scatter and intrinsic scatter is $\sim$ 0.02, and our estimates for $\beta$, $\sigma_{\rm BH}$, and $\sigma_{\star}$ are $\sim$ 1, 0.01 and 0.01, respectively. The physical sources of scatter from hierarchical merging and BH accretion are also shown in the upper panel of Figure 7. Note that at $z\sim 0.5$ the intrinsic scatter is less than the scatter due to hierarchical merging. This is because the scatter from different origins is obtained by independent numerical experiments that estimate approximate values for different terms in the simple model mentioned in Sect. 3.1, a procedure that may cause slight discrepancies in the final result due to the involved uncertainties.
Figure 7: Upper panel: The time evolution of the scatter of the $M_{\rm BH}-M_{\star}$ relation, the intrinsic scatter after correcting for the scatter from numerical origins, the scatter due to hierarchical merging, and the scatter due to BH accretion in the TNG100 simulation. Bottom panel: The time evolution of fractional contributions to the scatter due to numerical origins, hierarchical merging, and BH accretion in the TNG100 simulation. The statistical uncertainty of the results is $\sim 10^{-5}$. The figure indicates that the scatter from numerical origins increasingly dominates from $z=3$ to $z=0$, reaching $\sim$60% at $z=0$, whereas the scatter from accretion becomes less significant with time. At $z=0$, the scatter from physical origins is primarily contributed by hierarchical merging.
Note that the scatter from hierarchical merging is relatively low and increases with decreasing redshift, which arises because there massive galaxies increasingly merge with small galaxies that do not strictly lie on the $M_{\rm BH}-M_{\star}$ relation. In contrast, the inferred scatter from BH accretion decreases with decreasing redshift, and becomes consistent with zero at $z\lesssim 0.5$. This can be easily understood since the massive galaxies will tend to quench with decreasing redshift due to the AGN feedback, so that the contribution of BH accretion to the scatter scatter becomes lower and lower with time for massive galaxies and disappears at low redshift. We note that in realistic scenarios, mergers can indirectly enhance accretion by increasing $M_{\rm BH}$, which boosts the subsequent growth due to the $\dot{M}_{\rm BH}\propto M_{\rm BH}^{2}$ scaling. While this coupling is physically relevant, our approach treats such post-merger accretion as part of the accretion-driven component in order to maintain a clear separation between the direct contributions of hierarchical merging and black hole accretion.
@@
From the figure, it can also be seen that the scatter from the BH accretion decreases from a dominating $80\%$ at $z=3$ to something consistent with $\sim 0\%$ at $z\sim 1$. In contrast, the relative contribution of hierarchical merging increases from being negligible at $z=3$ to $\sim 40\%$ at $z\sim 0$, indicating that hierarchical merging contributes significantly to the scatter of the $M_{\rm BH}-M_{\star}$ relation at $z=0$ in the TNG100 simulations. It is however interesting to note that numerical sources are dominating the scatter at $z<2$ in the TNG model, therefore it would be misleading to directly compare the scatter obtained from the simulations with observations without a corresponding correction. In fact, the scatter seen in TNG100 at low redshift would be noticeable lower in the absence of numerical noise.
Figure 8: Upper panel: The time evolution of the intrinsic scatter of the $M_{\rm BH}-M_{\star}$ relation in TNG100, Illustris, EAGLE, and MTNG. Bottom panel: The time evolution of the fraction of the scatter due to BH accretion with respect to the total scatter in the TNG100, Illustris, EAGLE, and MTNG simulations, based on the assumption that the scatter from hierarchical merging is similar in all simulations. For the fraction of scatter from accretion, the four simulations show similar values at high redshift, but at $z=0$, a clear difference emerges, with MTNG and TNG showing a significantly lower proportion of scatter from accretion. This suggests that at low redshift the growth of black holes in MTNG and TNG is primarily driven by mergers.
Figure 9: Top panel: The time evolution of the total and intrinsic scatter in the $M_{\rm BH}-M_{\star}$ relation in the TNG100 and SeedVar simulations. The “SeedVar=0.5” simulation imposes a random scatter of 0.5 dex on the BH seed mass in a small TNG-like simulation with box size 25 $h^{-1}$ cMpc and $2\times 256^{3}$ particles. Middle panel: Scatter of BH mass across different mass bins in the SeedVar=0.5 simulation, stacked from $z=3$ to $z=0$. The shaded grey region denotes the 10-th and 90-th percentiles. Bottom panel: The time evolution of fractional contributions of different sources of scatter in the BH masses at different redshifts in SeedVar=0.5 simulation. The figure shows that the memory of an initial scatter in the BH seeds will be gradually erased and eventually becomes negligible in the TNG model due to the effective scatter reduction through mergers. We use a logarithmic $y$-axis scaling in the top and bottom panels. Since the values fall below $10^{-1}$ in some regions, this makes the differences between the lines easier to identify.
@@ 3.4 Other cosmological simulations and observational comparison
We now try to apply the lessons learned from TNG about the different sources of scatter in the $M_{\rm BH}-M_{\star}$ relation to the Illustris, EAGLE, and MTNG runs. Instead of carrying out similar numerical experiments as described in Section 3.2, we simply assume that the scatter due to hierarchical merging is similar in these simulations compared to TNG. This should be a particularly good assumption for the MTNG simulation, which shares the simulation code and an almost identical galaxy formation model with TNG100. Although we do not explicitly compare the BH merger rates between MTNG and TNG, the expected differences in merger-driven scatter due to cosmic variance or volume effects should be small, given their shared cosmological model and the statistically representative population of massive halos in both volumes. It is somewhat less obvious that this assumption applies to the Illustris and EAGLE simulations. However, their BH seeding scheme is conceptionally very similar to TNG, suggesting that the time evolution of the scatter due to hierarchical merging should be broadly similar.
@@
However, because the BH feedback and galaxy formation models in Illustris and EAGLE are quite different from TNG, the scatter due to numerical sources may be different. Thus we cannot directly apply the numerical scatter we obtained from TNG to the EAGLE and Illustris simulation data. Fortunately, Borrow et al. (2023b) measured the scatter due to the numerical variations in the EAGLE simulation. Based on their work, the variance of the BH mass $\sigma_{\rm BH}^{2}$ due to numerical origins in massive galaxies is around 0.010.030.01-0.03, and for the stellar mass the scatter $\sigma_{\star}^{2}$ in massive galaxies is around 0.01 from numerical origins, which is comparable to what we inferred for TNG. Unfortunately, for Illustris there is no work in the literature that has reported the scatter from numerical origins. As Illustris shares many aspects of the galaxy formation model with TNG, and also uses the same code, we however expect these two simulations to be similar in practice with respect to the numerical scatter, so we shall assume the TNG numerical scatter also for Illustris.
@@
Based on above assumptions, we can obtain the intrinsic scatter and the fractional contributions of different sources of scatter in these four cosmological simulations. We show the results in Figure 8, where the upper panel displays the intrinsic scatter in the four simulations, while the black dash line at $z\sim 0$ is the intrinsic scatter obtained by Kormendy & Ho (2013), who report a value of $\sim 0.29$ dex. This observational value is derived after correcting for typical measurement errors in BH mass estimates of $\sim 0.12$ dex. It can be seen that TNG100 and MTNG have very similar intrinsic scatter, as expected, which amounts to $\sim 0.1$ dex at $z\sim 0$. Note that the intrinsic scatter can appear slightly negative in MTNG in some cases. This arises because the numerical contribution is not measured within MTNG itself, but instead approximated using results from our controlled numerical experiments as mentioned before, which may not perfectly align with the simulation-specific noise properties. In contrast, Illustris and EAGLE show similar scatter, with values comparable to the observations. Since hierarchical merging reduces the scatter (Peng, 2007; Hirschmann et al., 2010; Jahnke & Macciò, 2011), and a tighter co-evolution of BHs and their host galaxies also regulates systems to be close to the $M_{\rm BH}-M_{\star}$ relation (Kormendy & Ho, 2013; Heckman & Best, 2014; Sun et al., 2015; Zhuang & Ho, 2023), we conclude that the low level of scatter in TNG100 and MTNG reflects a more efficient quenching of massive galaxies and a particularly tight self-regulation of the $M_{\rm BH}-M_{\star}$ relation in this model.
@@
The bottom panel of Figure 8 shows the fractional contribution of the variance from BH accretion to the total variance. This contribution by BH accretion is $50\%$ for TNG100 and MTNG at $z=3$ and then becomes negligible at low redshift. In contrast, in Illustris and EAGLE, the fraction of the variance from BH accretion is $\sim 80\%$ across the redshift range $z=3$ to $z=0$. This result indicates that the $M_{\rm BH}-M_{\star}$ relation is mainly regulated by hierarchical merging in TNG100 and MTNG at low redshift, with BH accretion becoming subdominant there, whereas Illustris and EAGLE appear to be less strongly quenched, so that variations in the BH mass by gas accretion are still more important. We will further discuss this issue in Section 4.2.
@@ 3.5 Scatter due to BH seed mass variations
As mentioned in Section 3.1, while the BH seed mass is fixed in the simulations, in the real Universe it may plausibly exhibit substantial variations, which could in turn affect the scatter in the $M_{\rm BH}-M_{\star}$ relation. The seed mechanism itself is presently unknown, and many different possibilities are actively discussed, including the run-away collapse of dense star clusters, the collapse of massive gas clouds, or the remnants of unusually massive pop-III stars. While the latter would still be comparatively light, the former may perhaps produce seeds with masses in excess of $10^{5}\,{\rm M}_{\odot}$. The seed masses could exhibit substantial variations (e.g. Bhowmick et al., 2024b), and especially if they are relatively massive the initial spread may still affect the scatter of the BH masses later on.
@@
Providing a more elaborate physical model for BH seeding, such as those based on local gas conditions and dynamical criteria (e.g. Tremmel et al., 2017), is beyond the scope of the current work. Instead we carry out a numerical experiment where we investigate the impact of varying the BH seed mass by running an extra cosmological simulation where we impose a random scatter of $0.5$ dex on the BH seed mass introduced in a small TNG-like simulation (hereafter SeedVar) with box size $25\,h^{-1}{\rm cMpc}$ and $2\times 256^{3}$ resolution elements.
@@
Figure 10: The scatter in the BH mass of galaxies of different stellar mass at different redshifts in the TNG100, Illustris, EAGLE, and MTNG simulations. The black line represents the scatters of BH mass for all galaxies. The blue line gives the scatter only for blue galaxies, while the red line correspondingly is for red galaxies. For the TNG100, Illustris, and MTNG simulations, the blue galaxies are defines as galaxies with color index (gr)<0.6(g-r)<0.6 while the red galaxies are the galaxies with color (gr)>0.7(g-r)>0.7. For EAGLE simulation, the blue galaxies are defined as galaxies with color (gr)<0.75(g-r)<0.75 while the red galaxies are those with color (gr)>0.75(g-r)>0.75. For TNG100, Illustris, and MTNG, there is a color gap since we ignore the green valley galaxies. Upper-left panel: TNG100 simulation. Upper-right panel: Illustris simulation. Lower-left panel: EAGLE simulation. Lower-right panel: MTNG simulation. The slightly higher scatter in blue galaxies than in red galaxies at $z=0$ indicates that the scatter contains the information about the growth history of BH. The consistent decrease of scatter with stellar mass at $M_{\star}\gtrsim 10^{10.5}\,{\rm M_{\odot}}$ at $z=0$, observed for both red and blue galaxies across all four simulations, suggests that this trend is driven by mergers.
@@
Figure 11: Left panels: The median value of the specific SFR as a function of stellar mass at different redshifts in the TNG100, Illustris, EAGLE, and MTNG simulations. Right panels: The median value of the BHAR as a function of stellar mass at different redshifts in TNG100, Illustris, EAGLE, and MTNG. The significant drop of the sSFR and the BHAR at the high mass end in TNG100/MTNG indicates that the accretion/in-situ star formation become less important for massive galaxies compared to Illustris/EAGLE.
@@
The result is presented in Figure 9. The top panel displays the evolution of variance and intrinsic scatter of the BH mass for the massive galaxies in the TNG100 and SeedVar simulations. The intrinsic variance is calculated by assuming that the numerical scatter in SeedVar is similar to that in TNG100. Comparing the resulting intrinsic variance in the SeedVar and TNG100 simulations, the scatter is higher in SeedVar than in TNG100 except for $z\sim 0$, where the intrinsic variance in SeedVar is comparable to that in TNG100. The variance in the SeedVar simulation at $z\lesssim 0.5$ experiences a steep decrease, from $\sim 0.1$ to $\sim 0.01$, i.e. the intrinsic scatter drops from $\sim 0.3$ dex to $\sim 0.1$ dex. This result is likely due to the effective galaxy quenching in TNG simulations.
@@
We further investigate the scatter of BH masses in different BH mass bins in the middle panel of Figure 9. We found that at different redshifts, the dependence of the scatter of BH masses in different BH mass bins is similar, so we stack the results from $z=3$ to $z=0$. The BHs in SeedVar show high scatter in low-mass BH bins due to the large variation of BH seeds. However, this increase in the scatter quickly decays away at $M_{\rm BH}\sim 10^{8}\,{\rm M_{\odot}}$, which can be interpreted as a loss of memory of the BH seed variations, and becomes $\lesssim 0.1\,{\rm dex}$ for massive BHs. This critical BH mass is the mass at which the impact of AGN feedback becomes significant, and the galaxies begin to be quenched (Zinger et al., 2020). At this stage, BHs also enter a rapid growth phase, during which their accretion is strongly modulated by AGN feedback. This self-regulated growth helps to suppress the divergence caused by initial seed differences and contributes to the flattening of the scatter distribution (see e.g. McAlpine et al., 2018). Since the quenched galaxies will have relatively low accretion rates due to a lack of gas supply, and the growth of BHs in quenched galaxies is mainly occurring due to mergers, we can infer that the continued drop of the scatter of the BH mass at high masses is primarily due to galaxy quenching and merger-driven averaging at $M_{\rm BH}\sim 10^{8}\,{\rm M_{\odot}}$ and above.
This transition mass marks the point where AGN feedback becomes significant and galaxies begin to quench (Zinger et al., 2020). At the same time, BHs enter a rapid growth phase in which accretion is strongly modulated by feedback, leading to self-regulated evolution. This process helps suppress the divergence caused by initial seed differences and contributes to the flattening of the scatter distribution (see e.g. McAlpine et al., 2018).
@@
Figure 12: Upper panel: The difference of the sSFR for the stellar mass bins in the ranges $10^{10-11}\,{\rm M_{\odot}}$ and $10^{10.5-11.5}\,{\rm M_{\odot}}$ in TNG100, Illustris, EAGLE, and MTNG. Lower panel: The difference of the BHAR in units of the Eddington rate for the stellar mass bins in the ranges $10^{10-11}\,{\rm M_{\odot}}$ and $10^{10.5-11.5}\,{\rm M_{\odot}}$ in TNG100, Illustris, EAGLE, and MTNG. Similar to Figure 11, the high difference of sSFR and BHAR in these figures also indicates that the growth of BH mass and stellar mass of massive galaxies proceeds in TNG100/MTNG mainly through dry mergers instead of in-situ star formation or BH accretion.